Added Sugars in a Typical Serving (in grams) Added Sugars in a Typical Serving (in grams) Foods Foods Table sugar, I tsp Bagel, cinnamon-raisin Bread, mixed-grain, slice Cake, chocolate, from pudding-type dry mix, slice Cake, snack-type, cream-filled, I Brownie, I Animal crackers, 12 Doughnut, cake-type, chocolate, sugared, or glazed Muffin, blueberry Pie, pecan, slice Pie, pumpkin, slice Toaster pastry, fruit Cola drink, 12 oz Ginger ale, 12 oz Note: Added sugars include sugar, fructose, lactose, maltose, glucose/dextrose, syrups, honey, molasses, and fruit juice concentrates. Chocolate syrup, 2 tbsp Citrus fruit drink, from frozen concentrate, 8 oz Cranberry juice cocktail, bottled, /2 c Corn flakes, I c Yogurt, fruit, nonfat, c Fast-food French toast sticks Salad dressing, Cae Applesauce, canned, sweetened% c Fruit cocktail, canned in heavy syrup, c Peanut butter, smooth, low-fat, tbsp Breakfast bar with oats, sugar, raisins Granola bar with oats, fruit, and nuts Candy, hard, I piece Chocolate pudding, ready-to-eat, /2 c 4 43 19 28 4 sar, low-calorie, 2 tbsp l0 34 38 32

A person consuming 1,800 kcal/day who is trying to limit added sugars to 10% of caloric intake would have a “budget” of about 180-sugar kcal (45 g) each day.

Look at Table 8-2 and see how difficult this might be. (above)

Assuming that you want to limit intake of added sugars to 15% of calories, and that you were consuming 1,800–2,600 kcal per day, calculate grams of added sugars allowable daily:

What foods would you need to give up? (Provide me with a minimum x6 examples)

What do you think that you would substitute for those foods?
Answer #1

1. Percentage of added sugar reduction is 15%

Total calories consumed were 1800-2600kcal

Then added sugar =15% of total calories 1800 kcal (taking the lower limit)

=15/100 × 1800

=15×18= 270 kcal of sugar

4kcal is present in 1gm of sugar

1kcal is present in 1/4gm of sugar

Then 270 kcal will be present in 1/4×270

= 67.5 gm of added sugar will provide 270kcal

Now taking the upper limit of calories 2600

Then added sugar =15%of 2600 = 390 kcal

4kcal is present in 1gm of sugar

1kcal is present in 1/4gm of sugar

Then 390 will be present in 1/4×390 = 97.5 gm

97.5 gm added sugar will provide 390kcal

2 The added sugars to be avoided are :-

- cake, chocolate, from pudding type dry mix,slice

- pie, pecan ,slice

- cola drink 12oz

- citrus fruit drink,from frozen concentrate

- yogurt, fruit ,nonfat

-fast food french toast sticks

3.Substitute for those foods are :-

- flavourful ,low calorie drink

- Greens like spinach, regular salad

- raw, cooked or roasted vegetables

- melon ,berries

- high fibres

- whole grains and protein